# On local Galois deformation rings: generalised reductive groups

### v1

@@ Appendix A Condensed sets
A condensed set can be defined as an accessible functor $X:\mathrm{ProFin}^{\mathrm{op}}\to\operatorname{Set}$ on the opposite category of profinite sets [48, Footnote 5], which satisfies the following two conditions:
@@
A map of condensed sets is a natural transformation. The category of condensed sets can be identified with the category of functors on the full subcategory of extremally disconnected profinite sets, which only satisfy condition (1) [48, Proposition 2.7]. We will switch between the two perspectives without further comment.
@@
For a surjection $f:X\to Y$ of condensed sets, we need to see that $f$ is the coequaliser of its kernel pair $X\times_{Y}X\rightrightarrows X$. This is a reflexive coequaliser and therefore sifted. The condensed sets $X$ and $Y$ can be seen as finite product preserving functors on the opposite category of the category of extremally disconnected profinite sets. Since sifted coequalisers commute with finite products, we can reduce the claim to sets. Every surjection of sets is an effective epimorphism, this completes the proof. ∎
@@
Suppose $X$ is a quasi-compact condensed set. By definition it arises as a left Kan extension from a $\kappa$-condensed set. By the density theorem we can write $X$ as a small colimit of $\kappa$-small profinite sets. The inclusion functor from $\kappa$-condensed sets into all condensed set preserves colimits [48, Proposition 2.9], so $X$ is a small colimit of profinite sets. In particular, there is an epimorphism $\coprod_{i\in I}\underline{S_{i}}\twoheadrightarrow X$, where $S_{i}$ are profinite sets. Since this epimorphism is effective by A.2 and $X$ is assumed quasi-compact, there is a finite subset $I_{0}\subseteq I$, such that $\coprod_{i\in I_{0}}\underline{S_{i}}\twoheadrightarrow X$ is a surjection.
Suppose there is a profinite set and a surjection $\underline{T}\twoheadrightarrow X$ and let $\coprod_{i\in I}X_{i}\twoheadrightarrow X$ be an effective epimorphism. By pullback we obtain a surjection $\coprod_{i\in I}(X_{i}\times_{X}T)\twoheadrightarrow T$. Since by [47, Proposition 1.2 (3)] $T$ is a quasi-compact object in the category of condensed sets, there is a finite subset $I_{0}\subseteq I$, such that $\coprod_{i\in I_{0}}(X_{i}\times_{X}T)\twoheadrightarrow T$ is an effective epimorphism. The map $\coprod_{i\in I_{0}}(X_{i}\times_{X}T)\twoheadrightarrow X$ factors over $\coprod_{i\in I_{0}}X_{i}\twoheadrightarrow X$, which is thus also a surjection. ∎
@@
Since $\underline{S}\times\underline{S}=\underline{S\times S}$ comes from a profinite set it is quasi-separated by [47, Proposition 1.2 (3)]. So $\underline{S}_{X\times X}X$ is quasi-compact.
@@
${T\times_{Y\times Y}Y}$${Y}$${X}$${T}$${Y\times Y}$${X\times X}$$\scriptstyle{\Delta_{Y}}$$\scriptstyle{\Delta_{X}}$
@@
By [47, Proposition 1.2 (4)] $X$ is a filtered colimit of quasi-compact quasi-separated condensed sets $X_{i}$ along injections. Recall, that the $X_{i}$ can be seen as compact Hausdorff spaces [47, Proposition 1.2 (3)]. So for any profinite set $S$

### v2

@@ Appendix A Condensed sets
A condensed set can be defined as an accessible functor $X:\mathrm{ProFin}^{\mathrm{op}}\to\operatorname{Set}$ on the opposite category of profinite sets [53, Footnote 5], which satisfies the following two conditions:
@@
A map of condensed sets is a natural transformation. The category of condensed sets can be identified with the category of functors on the full subcategory of extremally disconnected profinite sets, which only satisfy condition (1) [53, Proposition 2.7]. We will switch between the two perspectives without further comment.
@@
For a surjection $f:X\to Y$ of condensed sets, we need to see that $f$ is the coequaliser of its kernel pair $X\times_{Y}X\rightrightarrows X$. This is a reflexive coequaliser and therefore sifted. The condensed sets $X$ and $Y$ can be seen as finite product preserving functors on the opposite category of the category of extremally disconnected profinite sets. Since sifted coequalisers commute with finite products, we can reduce the claim to sets. Every surjection of sets is an effective epimorphism. This completes the proof. ∎
@@
Suppose $X$ is a quasi-compact condensed set. By definition it arises as a left Kan extension from a $\kappa$-condensed set. By the density theorem [41, Theorem III.7.1] we can write $X$ as a small colimit of $\kappa$-small profinite sets. The inclusion functor from $\kappa$-condensed sets into all condensed set preserves colimits [53, Proposition 2.9], so $X$ is a small colimit of profinite sets. In particular, there is an epimorphism $\coprod_{i\in I}\underline{S_{i}}\twoheadrightarrow X$, where $S_{i}$ are profinite sets. Since this epimorphism is effective by A.2 and $X$ is assumed quasi-compact, there is a finite subset $I_{0}\subseteq I$, such that $\coprod_{i\in I_{0}}\underline{S_{i}}\twoheadrightarrow X$ is a surjection.
Suppose there is a profinite set and a surjection $\underline{T}\twoheadrightarrow X$ and let $\coprod_{i\in I}X_{i}\twoheadrightarrow X$ be an effective epimorphism. By pullback we obtain a surjection $\coprod_{i\in I}(X_{i}\times_{X}T)\twoheadrightarrow T$. Since by [52, Proposition 1.2 (3)] $T$ is a quasi-compact object in the category of condensed sets, there is a finite subset $I_{0}\subseteq I$, such that $\coprod_{i\in I_{0}}(X_{i}\times_{X}T)\twoheadrightarrow T$ is an effective epimorphism. The map $\coprod_{i\in I_{0}}(X_{i}\times_{X}T)\twoheadrightarrow X$ factors over $\coprod_{i\in I_{0}}X_{i}\twoheadrightarrow X$, which is thus also a surjection. ∎
@@
Since $\underline{S}\times\underline{S}=\underline{S\times S}$ comes from a profinite set it is quasi-separated by [52, Proposition 1.2 (3)]. So $\underline{S}_{X\times X}X$ is quasi-compact.
@@
[…]
@@
By [52, Proposition 1.2 (4)] $X$ is a filtered colimit of quasi-compact quasi-separated condensed sets $X_{i}$ along injections. Recall, that the $X_{i}$ can be seen as compact Hausdorff spaces [52, Proposition 1.2 (3)]. So for any profinite set $S$
@@ A.1. Condensed structure on points
Glossary
$\operatorname{ad}\overline{\rho}$ adjoint representation associated to $\overline{\rho}$
$\operatorname{ad}^{0,\varphi}$ 3.6
$A^{\mathrm{gen}(,\tau)}_{G}$ coordinate ring representing $X^{\mathrm{gen},\tau}_{G}$
$A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ abbreviation of $A^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ as an $R^{\mathrm{ps}}_{G}$-algebra used after 8.4
$A^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ coordinate ring of $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$
$\mathfrak{A}_{\mathcal{O}}$ category of local Artinian $\mathcal{O}$-algebras with residue field $k$
$\mathfrak{A}_{\Lambda}$ category of local Artinian $\Lambda$-algebras with residue field $\kappa$
$A^{\chi}$ Section 15.2
$^{C}H$ $C$-group attached to a connected reductive group $H$
$\mathrm{CH}(D)$ Cayley–Hamilton ideal of a determinant law $D$
$\mathrm{cPC}^{\Gamma}_{G}(A)$ set of continuous $G$-pseudocharacters of $\Gamma$ with values in AA
$c_{P,L}$ semisimplification map $P(\kappa)\to G(\kappa)$, Section 2.2
$(-)_{\operatorname{disc}}$ discrete condensed set associated to a set
$\overline{D}$ determinant law associated to $\overline{\rho}$
$D_{\rho}$ determinant law associated to $\rho$
$D^{\square}_{(\rho,)G}$ framed deformation functor of $\rho$ with values in $G$, Section 3
$D^{u}$ universal determinant law extended to $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket$
$D^{u}_{|A}$ specialisation of $D^{u}$ along $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\to A$
$D^{u}_{E^{u}}$ determinant law induced by $D^{u}$ on $E^{u}$
$D_{\overline{\Theta}}$ deformation functor of $\overline{\Theta}$, 7.2
$D^{\square,\psi_{1}}_{G,\rho}$ Section 15
$E$ a finite extension of $F$ such that $\Gamma_{E}=\ker(\Gamma_{F}\xrightarrow{\overline{\rho}}G(k)\to(G/G^{0})(k))$
$E^{u}$ universal Cayley–Hamilton quotient $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket/\mathrm{CH}(D^{u})$
$\mathcal{F}_{N}$ free abstract group on $N$ generators
$G$ a generalised reductive group scheme over $\mathcal{O}$
$G^{\prime}$ derived group scheme of $G^{0}$
$G^{\prime}_{\operatorname{sc}}$ simply connected central cover of $G^{\prime}$
$G^{0}$ neutral component of $G$
$L_{\lambda}$ R-Levi attached to a cocharacter $\lambda$, Section 2.1
$^{L}H$ $L$-group of $H$, Section 16
$\mathfrak{m}_{A}$ maximal ideal of a local ring $A$, Section 3
NN fixed number of topological generators of $Q$, Section 5.1
$\mathcal{O}_{\kappa}$ Section 3
$\mathrm{PC}^{\Gamma}_{G}$ scheme of $G$-pseudocharacters of $\Gamma$, Section 7
$P_{\lambda}$ R-parabolic attached to a cocharacter $\lambda$, Section 2.1
$Q$ fixed quotient of $\Gamma$ as in 5.5
$\operatorname{Rep}^{\Gamma(,\square)}_{G}$ scheme of group homomorphisms $\Gamma\to G(A)$, Sections 1.4, 5.1
$R^{\mathrm{git},\tau}_{G}$ invariant ring $(A^{\mathrm{gen},\tau}_{G})^{G^{0}}$, coordinate ring of $X^{\mathrm{git},\tau}_{G}$, Section 6.1
$R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ abbreviation of $R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ as an $R^{\mathrm{ps}}_{G}$-algebra used after 8.4
$R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ connected component of $R^{\mathrm{git},\tau}_{G}$ corresponding to $\overline{\rho}^{\mathrm{ss}}$, Section 6.1
$R^{\square}_{(\rho,)G}$ universal deformation ring representing $D^{\square}_{\rho,G}$, Section 3
$R^{\square}_{\overline{\psi}}$ the universal framed deformation ring of $\overline{\psi}$
$R^{\square,\psi_{1}}_{\overline{\rho}}$ the universal deformation ring pro-representing $D^{\square,\psi_{1}}_{\overline{\rho}}$
$R^{\mathrm{ps}}_{G}$ the complete local noetherian $\mathcal{O}$-algebra pro-representing $D_{\overline{\Theta}}$, Section 7.2
$R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ universal deformation ring of $\overline{D}$, Section 5.1
$R^{\square,\psi_{1}}_{G,\rho}$ Section 15
$\underline{S}$ condensed set associated to a topological space $S$
$U^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ punctured space $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus\{\text{closed point}\}$
$U_{LG}$, $\overline{U}_{LG}$ Section 9.2
$U_{LG,\overline{\rho}^{\mathrm{ss}}}$, $\overline{U}_{LG,\overline{\rho}^{\mathrm{ss}}}$ component of $U_{LG}$, $\overline{U}_{LG}$ associated to $\overline{\rho}^{\mathrm{ss}}$
$U^{\mathrm{ps}}_{G}$ punctured spectrum $(\operatorname{Spec}R^{\mathrm{ps}}_{G})\setminus\{\text{closed point}\}$
$U_{\lambda}$ unipotent radical of $P_{\lambda}$, Section 2.1
$V_{LG}$, $\overline{V}_{LG}$ Section 9.2
$V^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$, $\overline{V}^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ 13.19
$\overline{V}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ 13.18
$X\sslash H$ affine GIT quotient of affine scheme $X$ by $H$, i.e. $\operatorname{Spec}(\mathcal{O}(X)^{H})$
$X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ Wang-Erickson’s space of generic matrices, 5.3
$X^{\mathrm{gen}(,\tau)}_{G}$ closed subscheme of $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ of representations factoring through $\tau$, 5.7
$\overline{X}^{\mathrm{gen},\tau}_{G}$ special fibre of $X^{\mathrm{gen},\tau}_{G}$
$X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ abbreviation of $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ as an $R^{\mathrm{ps}}_{G}$-scheme used after 8.4
$X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ component of $X^{\mathrm{gen},\tau}_{G}$ corresponding to $\overline{\rho}^{\mathrm{ss}}$, Section 6.2
$X^{\mathrm{gen},\psi_{1}}_{G,\overline{\rho}^{\mathrm{ss}}}$, $\overline{X}^{\mathrm{gen},\psi_{1}}_{G,\overline{\rho}^{\mathrm{ss}}}$ Section 15
$X^{\mathrm{gen}}_{P,\rho}$ 12.3
$X^{\mathrm{git}(,\tau)}_{G}$ GIT quotient $X^{\mathrm{gen},\tau}_{G}\sslash G^{0}$, Section 6
$\overline{X}^{\mathrm{git}(,\tau)}_{G}$ GIT quotient $\overline{X}^{\mathrm{gen},\tau}_{G}\sslash G^{0}$, Section 6
$X^{\mathrm{git}(,\tau)}_{HG}$, $\overline{X}^{\mathrm{git}(,\tau)}_{HG}$ image of $X^{\mathrm{git},\tau}_{H}\to X^{\mathrm{git},\tau}_{G}$ (resp. $\overline{X}^{\mathrm{git},\tau}_{H}\to\overline{X}^{\mathrm{git},\tau}_{G}$)
$X^{\mathrm{git}(,\tau)}_{HG,\overline{\rho}^{\mathrm{ss}}}$ component of $X^{\mathrm{git}(,\tau)}_{HG}$ corresponding to $\overline{\rho}^{\mathrm{ss}}$
$\overline{X}^{\mathrm{git}(,\tau)}_{HG}$ special fiber of $X^{\mathrm{git}(,\tau)}_{HG}$, as a topological space, 9.3
$X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ abbreviation of $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ as an $R^{\mathrm{ps}}_{G}$-scheme used after 8.4
$X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ connected component of $X^{\mathrm{git},\tau}_{G}$ corresponding to $\overline{\rho}^{\mathrm{ss}}$, Section 6.1
$X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ spectrum of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$, the universal deformation ring of $\overline{D}$, Section 5.1
$X^{\mathrm{ps}}_{G(,\overline{\rho}^{\mathrm{ss}})}$, $\overline{X}^{\mathrm{ps}}_{G(,\overline{\rho}^{\mathrm{ss}})}$ spectrum of $R^{\mathrm{ps}}_{G(,\overline{\rho}^{\mathrm{ss}})}$, $R^{\mathrm{ps}}_{G(,\overline{\rho}^{\mathrm{ss}})}/\varpi$
$X^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$, $\overline{X}^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$ 13.18
$X(\mu)$ group of characters $\mu\to\mathcal{O}^{\times}$
$X^{\chi}$ Section 15.2
$X^{\square}_{G,\overline{\rho}}$, $\overline{X}^{\square}_{G,\overline{\rho}}$ spectrum of $R^{\square}_{G,\overline{\rho}}$, $R^{\square}_{G,\overline{\rho}}/\varpi$
$Y$ preimage of closed points of $X^{\mathrm{git},\tau}_{G}$ in $X^{\mathrm{gen},\tau}_{G}$, Section 9.2
$Y_{\overline{\rho}^{\mathrm{ss}}}$ preimage of closed point of $X^{\mathrm{ps}}_{G}$ in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$
$Y^{\psi_{1}}_{\overline{\rho}^{\mathrm{ss}}}$ Section 15.2
$Z_{G}(H)$ centraliser of a subgroup scheme $H\leq G$
$Z(H)$ centre of a group scheme $H$, 2.1
$\underline{Z}(H)$ centre functor of a group scheme $H$, 2.1
$Z^{i}(\Gamma,V)$ set of continuous $i$-cocycles
$\overline{Z}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ 13.18
$\Delta$ image of $\pi_{G}\circ\overline{\rho}^{\mathrm{ss}}:\Gamma\to(G/G^{0})(\overline{k})$
$\underline{\Delta}$ constant group scheme associated to $\Delta$
$\Gamma_{F}^{\mathrm{ab},p}$ maximal abelian pro-$p$ quotient of $\Gamma_{F}$
$\gamma_{i}$ for $1\leq i\leq N$ fixed topological generators of $Q$, Section 5.1
$\overline{\Theta}$ the $G$-pseudocharacter attached to $\overline{\rho}$
$\Theta_{n}$ homomorphism $\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\to\mathrm{Map}(\Gamma^{n},A)$, part of $G$-pseudocharacter $\Theta$
$\Theta^{u}_{|A}$ specialisation of $\Theta^{u}$ along $R^{\mathrm{ps}}_{G}\to A$
$\Theta^{u}_{(G)}$ universal $G$-pseudocharacter over $R^{\mathrm{ps}}_{G}$
$\Theta_{\rho}$ $G$-pseudocharacter associated with $\rho$
$\kappa$ finite or local field in Section 3, otherwise a (topological) field
$\kappa(\mathfrak{p})$ residue field at prime ideal $\mathfrak{p}$
$\kappa[\varepsilon]$ ring of dual numbers $\kappa[\varepsilon]:=\kappa[x]/(x^{2})$ of $\kappa$
$\Lambda$ coefficient ring for $\kappa$, Section 3
$\Lambda^{0}$ subring of coefficient ring $\Lambda$ for $\kappa$, Section 3
$\mu_{p^{\infty}}(E)$ the group of all $p$-power roots of unity contained in $E$
$\pi_{1}(G^{\prime})$ kernel of $G^{\prime}_{\operatorname{sc}}\to G^{\prime}$
$\pi_{G}$ projection $G\to G/G^{0}$, Section 9
$\overline{\rho}$ a continuous representation $\Gamma_{F}\to G(k)$
$\rho_{A}$ a continuous representation $\Gamma_{F}\to G(A)$ lifting a given residual representation
$\rho^{\mathrm{ss}}$ $G$-semisimplification of $\rho$, 2.21
$\tau$ fixed faithful representation $\tau:G\to\operatorname{GL}_{d}$, Section 5.1
$\chi_{\mathrm{cyc}}$ the $p$-adic cyclotomic character
$\widehat{\otimes}$ completed tensor product, [59, Tag 0AMU]
$M^{\otimes_{R}d}$ abbreviation for the $d$-fold tensor product $M\otimes_{R}\cdots\otimes_{R}M$
$(-)^{\sharp}$ for a map $f:X\to Y$ of schemes $f^{\sharp}:\Gamma(Y,\mathcal{O}_{Y})\to\Gamma(X,\mathcal{O}_{X})$